OpenStudy (anonymous):

The circumference of a circle is 60π cm. What is the length of an arc of 140°?

OpenStudy (anonymous):

From Circumference given, first find radius.. \[2 \pi r = 60 \pi\]

OpenStudy (anonymous):

Calculate r here..

OpenStudy (anonymous):

still confused

OpenStudy (anonymous):

Where??

OpenStudy (anonymous):

can you tell me step by step?

OpenStudy (anonymous):

See, you are given with circumference.. Do you know the formula for Circumference??

OpenStudy (anonymous):

2 pi r

OpenStudy (anonymous):

or pi X diameter corrrect

OpenStudy (anonymous):

So just equate them and find r there..

OpenStudy (anonymous):

that's my problem... I don't kknow how to equate them

OpenStudy (anonymous):

\[C = 2 \pi r\] \[60 \pi = 2 \pi r\] Divide by 2pi both the sides to find r..

OpenStudy (anonymous):

I did it!! thank you. can I get more help?

OpenStudy (anonymous):

Sure.. If I know I will surely help..

OpenStudy (anonymous):

The corresponding sides of two similar rectangles are 6 ft and 10 ft. The area of the smaller rectangle is 78 ft2. To the nearest whole number, what is the area of the larger rectangle?

OpenStudy (anonymous):

Length of an Arc Formula= circumference * angle/360 =\[60\pi * 140/360\]

OpenStudy (anonymous):

i need new help

OpenStudy (anonymous):

so whats the answer? lol